# SlowPoke: Understanding and Detecting On-Chip Fail-Slow Failures in Many-Core Systems

### v1

@@ 3.2. SL-Compiler
Data map. Data collected by the probe fragments is stored in a performance data map, a highly configurable storage structure which balances precision and overhead. The map is parameterized by Level and Structure. The Level option controls aggregation granularity, from per-instruction (Inst) to per-stage (Stage), while the Structure option specifies storage organization. For example, we can use SL-Recorder (Sketch) for sparse data to save memory space and arrays (List) for dense data to provide fast access. To ensure low latency, these maps are initially stored in on-chip SRAM and are periodically offloaded to off-chip memory for long-running traces. Besides, by supporting incremental updates, eviction, and compression, these maps enable scalable trace management under different workload scales.
3.3. SL-Recorder
To reduce the storage overhead introduced by performance probes, we propose SL-Recorder, a lightweight on-the-fly data structure called the Fail-Slow Sketch, designed for trace compression and pattern-based filtering. As illustrated in Figure 6, the Fail-Slow Sketch operates in two stages. In Stage-1, a Running Track identifies frequently occurring traces with similar execution behaviors. In Stage-2, the common patterns among these traces are extracted, and their instance-specific information is aggregated based on shared patterns.
@@ 19 Function Insert_Stage_2($e$, attr)
Stage-1 consists of $d$ of bucket arrays, each associated with a distinct hash function. These hash functions map trace features that characterize execution patterns to bucket indices, associating traces with similar behaviors to the same bucket. Each array contains $m$ buckets, and each bucket has two fields: trace pattern (key) and its occurrence frequency. When the frequency exceeds a predefined threshold $H$, the corresponding trace pattern is considered a potential candidate for pattern extraction. Therefore, the threshold $H$ directly controls the similarity of trace patterns, while the total number of buckets influences accuracy indirectly by limiting the number of candidate traces tracked at the same time.
Stage-2 maintains a list of candidate fail-slow patterns and their associated attributes. This list supports operations such as updates and evictions, and its maximum length (MAX_LENGTH) is determined based on the characteristics of the workload. In this stage, traces with the same pattern are compressed, and statistical summaries of relevant performance metrics are computed to support root cause analysis. When the list reaches its maximum capacity, the Fail-Slow Sketch applies a First-In-First-Out (FIFO) replacement policy to make room for newly observed patterns.
@@
Update. Given a trace pattern and its attributes, the update operation modifies the corresponding entry in the Stage-2 pattern list to maintain the overall statistical information. For communication trace patterns, this includes updating metrics such as data volume and communication endpoints. For computation patterns, the update includes floating-point operation count, core id, and other relevant metrics. Additionally, all trace patterns have to maintain the timestamp of their first appearance and their duration, which are essential for localizing failure intervals during root cause analysis.
Figure 6. An example of Fail-Slow Sketch (d=1).
@@
Mathematical analysis. It can be proven that increasing the number of hash functions $d$ can effectively improve the retention probability, but also increases computational overhead; increasing the number of buckets $m$ helps reduce hash collisions, but also increases storage overhead. We provide a proof for the conclusion as follow.
@@
We further conduct parameter sensitivity analysis (Section 4.4) and design space exploration (Section 4.5) to determine the optimal configuration of SL-Recorder.
3.4. SL-Tracer
@@
SL-Tracer is designed to identify and localize fail-slow failures across both computation and communication resources. It follows a detection–construction–analysis pipeline. First, core-level detection and link-level detection are performed to identify potential fail-slow candidates at different stages of the system. These candidates, together with the workload’s communication dependency graph and the hardware topology, are integrated in the multi-level graph construction stage. The resulting graph captures both structural relationships and performance anomalies. Finally, the FailRank algorithm operates on this graph to quantify the likelihood of each candidate being the root cause, providing a ranked explanation of fail-slow components.
3.4.1. Core-Level Fail-Slow Detection
@@
The goal of core-level fail-slow detection is to identify candidate cores that exhibit abnormal slowdowns compared to their peers. Our method targets for structured parallel applications and performs this detection in a stage-aware manner, leveraging the mapping from workload to many-core architecture.
@@
As illustrated in Figure 7, we first partition computation instructions by execution stage (e.g., network layers, loop iterations, or pipeline segments), ensuring that only tasks with comparable semantics are grouped together. Next, cores are clustered according to the workload-to-core mapping. Each group therefore contains functionally equivalent computations executed across different cores, making their performance directly comparable.
@@
Within each group, we measure the performance of each core in terms of floating-point operations (FLOPs) per unit time. The group average serves as the baseline, and we perform outlier detection to highlight cores whose FLOPs are significantly lower than the baseline. These outliers are flagged as fail-slow candidates. For example, in Figure 7, core 6 and 10 deviate from their group peers and are thus marked as candidates. Each candidate is further assigned a fail-slow probability, derived from the normal distribution of its performance variance. This probability is then used as the initial fail-slow score for the FailRank algorithm (Sec 3.4.3).
To further localize fail-slow behavior, we maintain the start and end timestamps of each instruction. This temporal information reveals the intervals during which a slowdown exists, offering fine-grained visibility into the occurrence and persistence of fail-slow failure and enabling correlation between localized anomalies and downstream abnormal behaviors.
Figure 7. Core level fail-slow detection.
3.4.2. Link-Level Fail-Slow Detection
@@
The goal of the link-level fail-slow detection module is to identify interconnect links that exhibit abnormal bandwidth degradation during program execution. Since such fail-slow behaviors cannot be directly observed under limited hardware visibility, we infer them from communication traces collected by SL-Recorder. Specifically, we formulate the problem as a statistical inference task, where observed end-to-end communication delays are decomposed to estimate the bandwidth of individual links.
@@
For each communication event extracted from execution traces, we record four key attributes: the transferred data volume, the measured transmission time, and the index of the source and destination cores. Leveraging the deterministic routing algorithm of the 2D-mesh interconnect, we map each event to the set of links traversed along its multi-hop path, thereby associating observed end-to-end latency with individual link bandwidths.
@@
3.4.3. Fail-Slow Root Cause Analysis
To localize fail-slow root causes in many-core systems, we construct a multi-level communication graph that integrates hardware topology, execution dependencies, and runtime tracing data. This augmented graph captures how performance degradation propagates across cores and links, turning the localization into a graph-based inference problem. On this foundation, we design FailRank, a PageRank-inspired algorithm that quantifies the propagation path of fail-slows, enabling accurate root cause identification.
Multi-Level Communication Graph. We define the Multi-Level Communication Graph (MCG) as a layered graph $G=(V,E)$ that integrates execution dependency and hardware topology, shown in the right of Figure 8. Each node $v\in V$ represents a core at a specific time window, bounded with a set of computation instructions. The likelihood for being root cause of node $v$ is quantified by a fail-slow score $s(v)\in[0,1]$ (determined in Sec 3.4.1). Each directed edge $e=(u,v)\in E$ denotes communication dependency between two cores, bounded with a set of communication instructions. The edge is associated with a propagation weight $w(u,v)$, capturing the probability that performance degradation propagates from core $u$ to $v$.
Figure 8. Construction of multi-level communication graph.
@@ 3.4.3. Fail-Slow Root Cause Analysis
where $w(u,v)$ is the normalized propagation weight of edge $(u,v)$ defined in Sec 3.4.2. Each link’s score is updated after iteration by combining three factors: the propagation weight, the fail-slow scores of its endpoint node $u$, and its current score:
@@
3.5. Implementation
We implement SlowPoke simulator with SimPy(Müller et al., 2002), which enables event-driven modeling of concurrent events, resource contention, and time-dependent interactions across thousands of processes(Zhang et al., 2024). The many-core architecture, including cores, interconnects, and memory hierarchies, is represented as SimPy processes, allowing us to capture instruction traces, probe execution, and data movement for accurate modeling of fail-slow propagation. To evaluate failure scenarios, the simulator employs a monitoring-process-based injection mechanism: user-specified configurations are read at initialization, and during execution, a dedicated monitoring process dynamically alters the performance parameters of target components to emulate fail-slow behavior, automatically restoring them once the failure ends.
4. Evaluation
4.1. Experimental Setup
Evaluated Workloads. We construct a binary tree compute graph and calculate matrix operations on each node as a micro-benchmark. Besides, we evaluate SlowPoke on four representative neural networks, including GoogLeNet(Szegedy et al., 2015), DarkNet-19(Redmon and Farhadi, 2017), VGG(Simonyan and Zisserman, 2015), and ResNet-50(He et al., 2016). These workloads cover a range of computational and communication patterns, providing a comprehensive testbed for fail-slow detection. As mentioned in Sec 3.1, the workload mappings are generated using the Gemini(Cai et al., 2024) framework or following its style, which describes the tensor partition and block mapping clearly.
[…]
@@ 4.1. Experimental Setup
For different workloads, failures are reused as long as they can affect workload execution; failures that take place on unused resources are excluded. For different hardware architectures, additional failures are constructed to reflect the larger set of possible failure locations ($\times4$, $\times6$, $\times8$) in larger-scale systems. To ensure a balanced evaluation, we also construct an equal number of negative samples by modifying DNN structures to generate workloads without fail-slow failures. This dataset provides a controlled yet diverse set of fail-slow scenarios, enabling comprehensive evaluation of SLOTH in terms of detection accuracy, root cause localization, and robustness across workloads and architectures.
Baseline. To better evaluate the effectiveness of SlowPoke, we compare it with two baseline tracing approaches:
(1) Threshold Filtering: We implement a threshold based fail-slow detection method that directly analyzes the raw trace data collected during workload execution. This approach monitors key performance indicators such as computation time and communication latency, and flags a component as fail-slow if its metric exceeds a predefined threshold.
(2) Modified Microscope (Lin et al., 2018): This method uses directed acyclic graphs to represent the dependency among microservices, identifying and locating the abnormal services with a ranked list of possible root causes. We modified it to detect fail-slow failures.
4.2. Fail-slow Root Cause Detection
We evaluate SlowPoke’s effectiveness of fail-slow root cause detection across 5 workloads. As shown in Table 2, both Threshold Filtering and Microscop have low detection accuracy and high False-Positive Rate (FPR), indicating that they are not good at localizaing fail-slow root cause. In contrast, SlowPoke achieves consistently high accuracy, ranging from 80.4% to 94.98%. Among them, binary tree demonstrates the best performance, with 94.98% accuracy and a low FPR of 3.18%. The reason is the binary tree computation graph is structurally simple, making fail-slow attribution more reliable.
Table 2. Fail-slow detection accuracy.
[…]
Threshold DarkNet-19 50.92(138/271) 53.29(81/152)
GoogLeNet 51.65(141/273) 54.61(83/152)
VGG 45.45(125/275) 59.09(91/154)
ResNet-50 51.29(139/271) 56.77(88/155)
Binary Tree 60.44(165/273) 46.05(70/152)
Microscope DarkNet-19 68.20(193/283) 33.97(53/156)
GoogLeNet 67.14(188/280) 37.58(59/157)
VGG 68.46(191/279) 35.26(55/156)
ResNet-50 64.47(176/273) 45.22(71/157)
Binary Tree 79.20(217/274) 23.75(38/160)
SlowPoke (Ours) DarkNet-19 80.40(242/301) 12.64(22/174)
GoogLeNet 88.30(264/299) 13.64(24/176)
VGG 87.38(263/301) 14.61(26/178)
ResNet-50 82.79(255/308) 16.48(30/182)
Binary Tree 94.98(265/279) 3.18(5/157)
However, DarkNet-19 exhibits a relatively low accuracy (80.4%). The main reason lies in its uniform workload mapping pattern, which produces highly correlated traces. This correlation increases the estimation error of the EM-based inference, resulting in many undetected fail-slows. Other DNN workloads, such as GoogLeNet, VGG, and ResNet-50, fall between these two extremes, with accuracy around 83–88% and moderate error rates. Compared to DarkNet-19, these models benefit from more diverse communication patterns, which provide richer trace information for inference.
Overall, these results demonstrate that SlowPoke can achieve effective fail-slow root cause detection across different workloads, while its accuracy is influenced by workload structure and trace diversity.
4.3. Data Collection Overhead

### v2

@@ III-B SL-Compiler
Probe Abstraction. SL-Compiler decomposes the monitoring process into three configurable abstractions, as summarized in Table 1. Probe fragments define what data to collect—whether computation time (Exec), communication latency (Route), or memory access (Mem). Insertion tracepoints specify where to place probes by combining an instruction type filter (Type: Comp, Comm, or IO) with a relative position (Location: Pre, Post, or Surround). Data maps determine how to organize collected traces, supporting both fine-grained (Level: Inst) and coarse-grained (Level: Stage) aggregation, with flexible storage structures (Structure: List for dense access or Sketch for memory efficiency). Together, these abstractions enable users to express complex monitoring strategies through a compact five-tuple configuration.
Figure 5: Examples of SL-Compiler probing.
DNN-graph-guided Insertion. A key innovation of SL-Compiler is its ability to automatically generate probe configurations by analyzing the DNN computation graph and its mapping to the target architecture. As illustrated in Figure 5, this process takes as input the workload mapping—which specifies how DNN operators are distributed across cores—and performs static analysis to identify critical monitoring points. The process proceeds in three steps. First, it parses the computation graph to extract the layer sequence, data dependencies, and operator types (e.g., Conv2D, Pooling). Second, it analyzes the workload-to-core mapping to determine which cores execute communication-intensive operations (requiring Route probes) versus computation-heavy operators (requiring Exec probes). For instance, in Figure 5b, it identifies that Conv layers dominate computation and automatically inserts post-execution probes ([Exec, Comp, Post, Stage, List]) to capture their execution time. Similarly, in Figure 5c, it detects inter-core data transfers and inserts pre-communication probes to monitor link latency. Third, it generates the instrumented code by inserting lightweight probe fragments at the identified tracepoints. Importantly, this process is fully automated, freeing users from manual instrumentation while still supporting custom probe specifications for specialized analysis needs.
Operator-level Semantic Collection. Beyond basic performance metrics, SL-Compiler captures high-level DNN operator semantics to facilitate root cause analysis. Each probe records not only the standard execution attributes—such as start/end timestamps, and core ID, but also DNN-specific metadata that characterizes the computational context. For computation probes, this includes the operator type (e.g., Conv2D, GEMM), input/output tensor dimensions, and floating-point operation count (FLOPs). For communication probes, it records the data volume transferred, the source and destination cores, and the corresponding tensor being communicated. This semantic information serves two critical purposes: (1) it enables layer-aware analysis in SL-Tracer, allowing the framework to correlate slowdowns with specific DNN layers; and (2) it facilitates communication pattern recognition, helping distinguish between normal inter-layer data transfers and anomalous traffic caused by fail-slow failures. For example, if a Conv layer exhibits unexpectedly high communication latency, the recorded tensor dimensions can reveal whether the slowdown stems from excessive feature map sizes or from degraded link bandwidth.
III-C SL-Recorder
To reduce the storage overhead, we propose SL-Recorder, a lightweight on-the-fly data structure for trace compression and pattern-based filtering. The data structure is also called Fail-Slow Sketch as it is inspired by sketches. As illustrated in Figure 6, Fail-Slow Sketch operates in two stages. In Stage-1, a Running Track identifies frequently occurring traces with similar execution behaviors. In Stage-2, the common patterns of the traces are extracted to reduce storage redundancy.
@@ 19 Function Insert_Stage_2($e$, attr)
Stage-1 consists of $d$ bucket arrays, each associated with a distinct hash function. These hash functions map trace features that characterize execution patterns to bucket indices. Each array contains $m$ buckets, and each bucket has two fields: trace pattern (key) and its frequency. When the frequency exceeds a threshold $H$, the corresponding trace pattern is considered as a potential candidate for fail-slow pattern.
Stage-2 maintains a list of candidate fail-slow patterns and their associated attributes. This list supports operations such as updates and evictions, and its maximum length (MAX_LENGTH) is determined based on the characteristics of the workload. In this stage, traces with the same pattern are compressed, and statistical summaries are computed to support root cause analysis. When the list reaches its maximum capacity, the Fail-Slow Sketch evict the most “healthy” pattern to make room for newly observed patterns.
@@
Figure 6: An example of Fail-Slow Sketch (d=1).
Update. Given a trace pattern and its attributes, the update operation modifies the corresponding entry in the Stage-2 pattern list to maintain the overall statistical information. For communication patterns, this includes updating metrics like data volume. For computation patterns, the update includes floating-point operation count, core id, and other relevant metrics. Additionally, all trace patterns have to maintain the timestamp of their first appearance and their duration, which are essential for root cause analysis.
@@
Mathematical Analysis. It can be proved that increasing the number of hash functions $d$ can effectively improve the retention probability, but also increases computational overhead; increasing the number of buckets $m$ helps reduce hash collisions, but also increases storage overhead. We provide a proof for the conclusion as follow.
@@
We further conduct parameter sensitivity analysis (Section IV-D) and design space exploration (Section IV-E) to determine the optimal configuration of SL-Recorder.
III-D SL-Tracer
@@
SL-Tracer is designed to identify and localize the root cause fail-slow failures. It follows a detection–construction–analysis pipeline. First, core-level detection and link-level detection are performed to identify potential fail-slow candidates. These candidates, together with workload’s communication dependency graph and the hardware topology, are integrated together to construct multi-level communication graph (MCG). Finally, the FailRank algorithm operates on this graph to quantify the likelihood of each candidate being the root cause, providing a ranked explanation of fail-slow components.
III-D1 Core-Level Fail-Slow Detection
@@
The goal of core-level fail-slow detection is to identify candidate fail-slow cores. Our method performs the detection in a stage-aware manner, leveraging the mapping from DNN to many-core accelerator.
@@
As illustrated in Figure 7, we first partition computation instructions by execution stage (e.g., network layers), ensuring that only tasks with comparable semantics are grouped together. Next, cores are clustered according to the workload-to-core mapping. Each group therefore contains volume-equivalent computations executed across different cores, making their performance directly comparable.
@@
Within each group, we measure the performance of each core in terms of floating-point operations (FLOPs) per unit time. The group average serves as the baseline, and we perform outlier detection to highlight cores whose FLOPs are significantly lower than the baseline. These outliers are flagged as fail-slow candidates. For example, in Figure 7, core 6 and 10 deviate from their peers and are thus marked as candidates. Each candidate is further assigned a fail-slow probability, derived from its performance variance distribution. This probability is then used as the initial fail-slow score for the FailRank algorithm (Sec III-D3).
Figure 7: Core level fail-slow detection.
III-D2 Link-Level Fail-Slow Detection
@@
The goal of link-level fail-slow detection is to identify candidate fail-slow interconnect links. Since such fail-slow behaviors cannot be directly observed under limited hardware visibility, we infer them from communication traces collected by SL-Recorder. Specifically, we formulate the problem as a statistical inference task, where observed end-to-end communication delays are decomposed to estimate the bandwidth of individual links.
@@
For each communication event, we record four key attributes: the transferred data volume, the measured transmission time, and the index of the source and destination cores. Leveraging the deterministic routing algorithm of the 2D-mesh interconnect, we map each event to the set of links traversed along its multi-hop path, thereby associating observed end-to-end latency with individual link bandwidths.
@@
To assess the likelihood of fail-slow behavior, we calculate the fail-slow probability using the bandwidth gamma distribution model. These probabilities are then assigned as the initial fail-slow scores of links in the FailRank algorithm.
III-D3 Fail-Slow Root Cause Analysis
To localize fail-slow root causes, we construct MCG that integrates hardware topology, execution dependencies, and runtime tracing data. This augmented graph captures how performance degradation propagates across cores and links, turning the localization into a graph-based inference problem. On this foundation, we design FailRank, a PageRank-inspired algorithm that quantifies the propagation probability of fail-slows, enabling accurate root cause identification.
Figure 8: Construction of multi-level communication graph.
MCG Definition. MCG is defined as a layered graph $G=(V,E)$ that integrates task dependency and hardware topology, shown in the right of Figure 8. Each node $v\in V$ represents a core in a time window, bounded with a set of computation instructions. The likelihood for being root cause of node $v$ is quantified by a fail-slow score $s(v)\in[0,1]$ (determined in Sec III-D1). Each directed edge $e=(u,v)\in E$ denotes communication dependency between two cores, bounded with a set of communication instructions. The edge is associated with a propagation weight $w(u,v)$, capturing the probability that performance degradation propagates from core $u$ to $v$.
@@ III-D3 Fail-Slow Root Cause Analysis
where $w(u,v)$ is the normalized propagation weight of edge $(u,v)$ defined in Sec III-D3. Each link’s score is updated by three factors: the propagation weight, the fail-slow scores of its endpoint node $u$, and its current score:
@@
IV Evaluation
IV-A Experimental Setup
Evaluation Platform. We implement an event-driven simulator for many-core DNN accelerators using SimPy[52], a Python-based discrete-event simulation framework. The simulator can capture concurrent execution across thousands of processes, accurately modeling resource contention and inter-core communication. To ensure the reliability of our simulator, we validated it against the Dataflow Abstract Machine (DAM) framework[82], a parallel simulator for dataflow systems. We cross-verified total execution cycles produced by our simulator with DAM’s Communicating Sequential Processes with Time (CSPT) model on several dataflow graphs under identical workload configurations.
To faithfully capture real-world hardware behavior, our simulator incorporates statistical models[59, 29] for key performance parameters based on established distributions observed in production systems[32, 49]. For computing capacity, we model the computing capacity $P_{core}$ of each core as a normally distributed random variable: $P_{core}\sim N(\mu_{c},\sigma^{2}_{c})$ where $\mu_{c}$ represents the expected processing capacity and $\sigma^{2}_{c}$ quantifies the runtime variability. For link bandwidth, we model the transmission time $T_{link}$ using a Gamma distribution: $T_{link}\sim Gamma(\alpha,\beta)$ to capture latency variation caused by bandwidth. The shape parameter controls the asymmetry of latency distribution, while the rate parameter reflects the average link throughput.
TABLE III: Fail-slow detection accuracy.
Algorithm DarkNet-19 GoogLeNet VGG ResNet-50 Binary Tree
Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%) Acc.↑(%) FPR↓(%)
Thres 50.92(138/271) 53.29(81/152) 51.65(141/273) 54.61(83/152) 45.45(125/275) 59.09(91/154) 51.29(139/271) 56.77(88/155) 60.44(165/273) 46.05(70/152)
Mscope 68.20(193/283) 33.97(53/156) 67.14(188/280) 37.58(59/157) 68.46(191/279) 35.26(55/156) 64.47(176/273) 45.22(71/157) 79.20(217/274) 23.75(38/160)
IASO 31.65(163/515) 58.58(215/367) 37.14(156/420) 43.28(116/268) 22.53(164/728) 73.93(433/585) 13.54(221/1632) 89.85(1347/1499) 38.27(175/457) 52.20(166/318)
Perseus 66.38(235/354) 32.44(73/225) 80.75(235/291) 6.17(10/162) 45.37(235/518) 60.93(237/389) 16.84(235/1395) 87.99(1114/1266) 78.33(235/300) 11.11(19/171)
ADR 10.58(235/2222) 92.73(1941/2093) 11.17(235/2102) 92.29(1821/1973) 11.06(235/2123) 92.37(1842/1994) 11.17(235/2102) 92.29(1821/1973) 11.30(235/2079) 92.20(1798/1950)
SLOTH (Ours) 80.40(242/301) 12.64(22/174) 88.30(264/299) 13.64(24/176) 87.38(263/301) 14.61(26/178) 82.79(255/308) 16.48(30/182) 94.98(265/279) 3.18(5/157)
Workloads. We evaluate SLOTH on four representative DNN models that exhibit diverse characteristics: GoogLeNet[65], DarkNet-19[57], VGG[63], and ResNet-50[26]. These workloads span a range of architectural patterns from highly parallel branching structures to memory-intensive linear pipelines. The input dimensions of all workloads are configured to match typical throughput-oriented inference (batch size = 64). For controlled microbenchmark analysis, we additionally construct a binary tree computation graph where each node performs matrix operations. This synthetic workload isolates communication patterns from complex DNN semantics, facilitating ablation studies on probe latency and storage overhead.
The mapping configurations are generated by the Gemini framework[7], which specifies the tensor partitioning strategy and the block-to-core assignment. This deterministic mapping allows us to inject fail-slow failures at specific cores and observe their propagation through the computation graph.
@@ IV-A Experimental Setup
Fail-Slow Dataset. We construct a fail-slow dataset comprising 152 failure instances characterized by three parameters: failure location (core-level or link-level at a 7:3 ratio), duration (uniformly distributed between 0–10 s), and slowdown rate (fixed at 10×). Failures are reused across workloads if they affect execution; those occurring on unused resources are excluded. For larger architectures (4×4, 6×6, 8×8), we generate additional failures proportional to the expanded resource count. To ensure balanced evaluation, we construct an equal number of negative samples by modifying DNN structures to produce workload traces without fail-slow failures. This dataset enables comprehensive evaluation of detection accuracy, localization precision, and robustness across diverse failure scenarios, workloads, and system scales.
Baseline. We compare SLOTH against five baseline methods originally designed for microservices, distributed storage, or server clusters. We adapt these methods to the many-core DNN accelerator domain with in-house implementations, and all the methods share the same raw trace collection infrastructure for fair comparison. For methods providing risk-level outputs, we classify moderate risk or higher as fail-slow. The parameters used by each method have been searched to match our data. Detailed description of baselines is as follows:
(1) Threshold Filtering (Thres) monitors raw performance metrics (per-core computation time, link latency) and flags components exceeding 2× the nominal latency threshold derived from workload profiling as fail-slow.
(2) Microscope (Mscope) [40] constructs a directed acyclic graph of inter-service dependencies and ranks root causes using random walk-based scoring. We model data dependencies between operators as service dependencies.
(3) IASO [55]. A peer-based detection framework that converts timeout signals into scores using AIMD combined with DBSCAN clustering. We treat each core as a peer node, using execution time violations as timeout signals.
(4) Perseus [46] builds polynomial regression models on latency-versus-throughput distributions, employing PCA and DBSCAN for outlier detection. We construct core-level model and use the 99.9th percentile as the detection threshold.
(5) Adaptive Detection at Runtime (ADR) [45] monitors both value and update frequency of performance variables using sliding windows with adaptive thresholds based on historical percentiles.
IV-B Fail-slow Root Cause Detection
We evaluate detection effectiveness across five workloads. As shown in Table III, baseline methods exhibit distinct limitations that impact their accuracy.
Thres and Mscope can detect both node and link fail-slows but suffer from low sensitivity. Thres misses 40-50% of true fail-slows across workloads, as static thresholds cannot adapt to dynamic performance variations. Mscope achieves slightly better recall but still misses 20-32% of cases, and its FPR reaches 33-54% due to incorrect attribution in complex dependency graphs. IASO, Perseus, and ADR detect more true fail-slows but are limited to node-level detection. More critically, these methods misclassify propagated slowdowns as root causes, resulting in high FPR. ADR’s adaptive thresholding suffers from the highest FPR (91-93%) by treating persistent propagated effects as independent failures, while Perseus shows workload-dependent behavior: it achieves low FPR (6.17%) on structurally simple GoogLeNet but degrade significantly on complex models like ResNet-50 (FPR: 87.99%).
Workload characteristics significantly influence detection difficulty. Binary Tree’s simple structure and regular communication patterns yield the highest accuracy across all methods. Conversely, DarkNet-19’s uniform workload mapping produces highly correlated traces, degrading inference quality—even SLOTH achieves only 80.4% accuracy here. Complex CNNs (GoogLeNet, VGG, ResNet-50) fall in between, where diverse layers provide richer trace information but also introduce more propagation paths that confuse baseline methods.
SLOTH outperforms baselines by addressing both node and link fail-slows while explicitly modeling fault propagation through iterative inference. It achieves 80-95% accuracy with low FPR (3-17%), demonstrating robust performance across different workload structures.
Figure 10: Time overhead of SL-Compiler probing.
IV-C Data Collection Time Overhead